Math 242 (Section 3)  University of South Carolina
Prof. Meade  Spring 1998

Supplemental Course Evaluation

Please return these forms in class on Thursday, April 23. You do not have to give your name, but if you wish to amplify your comments in person, please feel welcome to do so. Constructive criticism is appreciated! Thank you.

1. State three mathematical things that you have learned in the class so far.

(a)

(b)

(c)

2. State three non-mathematical things that you have learned in the class so far.

(a)

(b)

(c)

3. Do you feel the applied aspects of the course (word problems and projects) have helped you learn? Would you like to see more applications? If so, do you have any specific suggestions?

4. My impression is that when most of you use Maple you don’t initiate the commands yourself, but rather you re-use bits and pieces of the Maple code that I have distributed, making modifications as necessary. Is this impression correct? Do you think this is an effective way to use the software? Is there anything that could (or should) be done to facilitate your use of Maple (in or out of class)?

5. Concerning the integrated use of Maple in this class: did you find this aspect of the course enjoyable? did you find it educational? was this a more effective way of learning Maple than watching me give a demonstration? would you have preferred more or less use of the computer?

6. Are you getting enough feedback from graded quizzes? Would you have preferred to have graded homework? Did you feel adequately prepared for the exams?

7. Concerning the text. How are the problem sets: useful for learning skills? for learning concepts? at the right level? interesting?

8. How were the supplemental worksheets and other handouts: useful for learning skills? for learning concepts? at the right level? interesting?

9. Did group work help you learn in this course? How effectively did your informal and formal groups function? Should the instructor take more control over the formation of groups for the projects?

10. Which aspects of the course would you like to change? to keep the same?

11. Have you enjoyed the class?

12. Would you recommend this class to a friend who will be taking this course in the future?

Computer/Maple survey from RPI

Please mark the appropriate box with an × or √.

Strongly Slightly  Strongly Slightly
Disagree Disagree Neutral Agree Agree

1. I enjoyed the course.

2. The computer sessions and assignments were interesting.

3. The computing environment was introduced satisfactorily.

4. The computer work was reasonable in length and difficulty.

5. Working on the computer with a partner was helpful and worthwhile.

6. Maple enables me to solve some problems that would be almost impossible by hand calculations.

7. By using Maple I am able to see new approaches for solving some problems.

8. Maple was easy to apply to a wide variety of problems.

9. The use of Maple in this course has improved my problem-solving skills.

10. The use of Maple revealed aspects of mathematics (and ODEs in particular) ODEs that I hadn’t thought about before.

11. The use of Maple was satisfactorily interwoven with the rest of the course.

12. My knowledge of Maple will probably help me in other science and engineering courses.

13. I would rather have had an ODE course that did not use Maple.

14. What was the most useful/enjoyable part about using Maple?

15. What was the least useful/enjoyable part about using Maple?

Project Questionnaire

Please respond to the following questions pertaining to the projects from this class. If you have additional comments, please feel free to attach additional sheets.

Strongly Slightly  Strongly Slightly
Disagree Disagree Neutral Agree Agree

1. The projects helped me gain confidence in my ability to solve scientific problems.

2. The projects helped me develop my groupwork skills.

3. The projects helped me develop my independent analytical thinking skills.

4. The projects helped me develop the ability to mathematically model a situation by defining variables, establishing relationships between them, drawing conclusions, and interpreting the results.

5. The projects helped me develop my technical writing skills.

6. The projects increased my appreciation of mathematics as a subject that is evolving, and not all finished and known.

7. I would rather have an ODE course that did not have projects.

8. What was least useful/enjoyable about the projects?

9. What was most useful/enjoyable about the projects?